

********************************************************************
UARS WEEKLY STATUS REPORT - 26 JAN 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE is in good health, collecting solar data on all available orbital
tracking periods and stellar calibration observations whenever possible. We
continue to gather enough data to meet all the SOLSTICE science objectives
during this period of reduced tape recorder operations.
***********
SUSIM
***********
SUSIM performed its daily solar scans of the sun during periods of
predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Assisted the FOT in retrieving data associated with past
spacecraft anomalies. Daily processing of instrument Daily
Activity Plans continues with no problems.
***********
CDHF
***********
STATUS:
Data created in the month of December, 2000 and for the first 17 days of
January, 2001 has been loaded into the DAAC Transfer Tables for MLS V5.
DAAC has been notified.
UARS CDHF REHOST:
- We are making plans to go operational on the Rehost System
31-JAN-2001.
- PEM V3.32 (JOBS 1 and 2) have both tested successfully on the
Rehost System. Currently waiting validation results from the PEM team.
***********
FOT
***********
REPORTING PERIOD
This report covers 13 January 2001 (Orbit 51061, GMT Day 013) through 19
January 2001 (Orbit 51165, GMT Day 019).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 013/0235 51062 Power OFF
The beta angle ranged between the angles of 55.8 degrees to 76.1 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 77 degrees on 21 January 2001 (DOY 021).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.3 and 0.9 degrees C. Additional battery performance data for each
battery on 12 January 2001 and 19 January 2001 are provided in Appendix D.
The clock error ranged from Ð12 to 12 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 05
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 09
¥ HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 03
Five (5) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/315 - 2287.496493
· Transponder B: 00/314 - 2287.498584
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/018 show the observatory data loss to be 300 hours, 08
minutes, 41.256 seconds. This shows an overall increase of 0 seconds since the
last report period This is a 0.39 percent data loss, which equals a 99.61
percent data capture for the mission.
· There was no data loss for the week.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UARS began flying through the high beta period on 16 January 2001, with the TAM
as its only input to UFLTR. The TAM is maintaining attitude control to within
science requirements as planned. More fine-tuning of the TAM is needed before
it will be operating at its maximum effectiveness, but it is at present
performing very well.
MEETINGS
The FOT attended a TAM Attitude Meeting on 18 January 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 09 February 2001 (DOY 040). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 19 January 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
51062 013/0235 01-042 HALOE Power OFF
51063 013/0354 01-043 OBC Clk Adj from 20.6 to 0.00 msec/day
51070 013/1442 01-044 OBC TAM ALIGNMENT
51072 013/1915 01-045 NBTR NOOPS D014
51080 014/0757 01-045 NBTR D-014 NOOPs
51087 014/1827 01-046 OBC Clk Adj from 0.00 to 20.6 msec/day
51118 016/2003 01-047 OBC Clk Adj from 20.6 to 41.6 msec/day
51118 016/2100 01-048 OBC MAGMOD Time Update
51123 017/0437 01-049 OBC ZEPS & NEPS on full time
51160 019/1601 01-050 OBC Clk Adj from 41.2 to 20.6 msec/day
__________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 19 January 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
QTZEPSIF2 Y-LO I/F Plate Forward 51119-20
QTZEPSIF1 Y-LO ZEPS Interface Temp 1 51121-35
QTZEPSIF2 Y-LO ZEPS Interface Temp 2 51121-35
NBTUTEMP2 Y-LO NBTR TU-2 TEMP 37 51155
NBTUTEMP2 R-LO NBTR TU-2 TEMP 39 51156
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 19 January 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
51063 013/0400 Clk Adj from 20.6 to 0.00 msec/day
51087 014/1836 Clk Adj from 0.00 to 20.6 msec/day
51118 016/2011 Clk Adj from 20.6 to 41.6 msec/day
51160 019/1622 Clk Adj from 41.2 to 20.6 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 19 January 2001
12 Jan 2001 - GMT Day 012 Beta = 51.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/+0.0 14.2 0.000 0.0
2 +11.2/-95.2 +5.15/+2.92 +25.6/-15.2 25.1 1.050 19.9
3 +22.4/-44.8 +3.87/+1.97 +25.6/-15.6 25.1 1.040 19.7
19 Jan 2001 - GMT Day 019 Beta = 76.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ------- --------- ----- ------- ------
1 +728.0/+728.0 -2.42/-3.68 +0.0/+0.0 14.2 0.000 0.0
2 +11.2/-22.4 +3.23/+1.66 +15.2/-16.0 27.4 1.050 14.1
3 +11.2/-28.0 +2.29/+1.34 +14.4/-15.6 27.4 1.040 13.8
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
---- ---- ---------
-no changes V/T 5
___________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 19 January 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-002 FOT 51076 014/0056 14170 Generic Late Acquisition
01-003 FOT 51111 016/0856 N/A AP Crashed/disk crash
01-002 FOT 51148 018/2005 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-053 FOT 50168 319/1113 23157 Late Acquisition
00-054 FOT 50280 326/2147 23187 RESET PTP /2:39 32K
00-056 FOT 50584 347/0502 23302 Premature LOS
01-002 FOT see list Jan 01 14170 Generic Late Acquisition
01-003 FOT 51111 016/0856 N/A AP Crashed/disk crash
SEVNTFW01/03